OpenStudy (crashonce):

Easier trig help pls sorry tan(A+60)tan(A-60)(1+tanA)=-3 @zepdrix @ikram002p @ganeshie8

OpenStudy (crashonce):

@ganeshie8 any ideas? just a formula to use will do if it works

ganeshie8 (ganeshie8):

the donkey way would be to apply tan(A+B) formula and try simplifying all day...

OpenStudy (crashonce):

haha yea i thought abt that but is there a better way?

OpenStudy (ikram002p):

apply soemthing for tan (a) tan (b ) maybe ?

OpenStudy (crashonce):

ye but wats the tanAtanB formula i still dont know lol

ganeshie8 (ganeshie8):

actually tan(a)tan(b) formula simplifies smoothly

ganeshie8 (ganeshie8):

\[\large \tan A \tan B = \dfrac{\tan A - \tan B}{\tan(A-B)} - 1\]

ganeshie8 (ganeshie8):

\[\tan(A+60)\tan(A-60)(1+\tan A)\] \[\left(\dfrac{\tan(A+60)+\tan(A-60)}{\tan(A+60-(A+60))}-1\right) (1+\tan A)\]

ganeshie8 (ganeshie8):

\[\left(\dfrac{\tan(A+60)+\tan(A-60)}{\tan 120}-1\right) (1+\tan A)\]

OpenStudy (midhun.madhu1987):